OpenStudy (anonymous):

What is the solution of the equation? 6x – 3 = 5x – 5

OpenStudy (zzr0ck3r):

6x-3 = 5x-5 gather like terms 6x-5x=-5+3 x = ?

OpenStudy (anonymous):

x=2 britbrat

OpenStudy (zzr0ck3r):

nope

OpenStudy (zzr0ck3r):

what is -5+3?

OpenStudy (anonymous):

-2 forgot the sign

OpenStudy (zzr0ck3r):

we dont just give answers here, there is not much point in that.

OpenStudy (lgbasallote):

i think it's their avenue of showing how "smart" they are @zzr0ck3r

OpenStudy (anonymous):

that u idepallo